• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name JPCHEN

    Log entry time 14:26:35 on October 4,1998

    Entry number 2315

    keyword=changed to longitudinal pumping

    Transverce completed for P8 at 2:10 pm. Turned off transverce lasers, rotated to longitudinal and did NMR. Polarization EPR/NMR=28.29/27.16 %. In Empty position now for BCM calibration.